OpenStudy (anonymous):

3x-12=18x please show work

OpenStudy (anonymous):

So I'm assuming you're trying to solve for x?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Well, first I would subtract 3x from 18x which would result in -12=15x. then you would divde by -12 on both sides. -12/-12=15x/-12 that would simplify into the fraction x= -5/4

OpenStudy (solomonzelman):

\[3x-12=18x\]\[3x-12\color{red} {-3x}=18x\color{red} {-3x}\]\[-12=15x\]\[-12\color{red} {\div (15)}=15x\color{red} {\div (15)}\]\[-12/15~~~=~~~x\]

OpenStudy (solomonzelman):

Good?

OpenStudy (anonymous):

Oh poop. I'm crazy @SolomonZelman is right. Sorry! I got a little muddled.

OpenStudy (anonymous):

technically.... you could factor it first to make life a bit easier :P but whatever works